A sample of nitrogen at 20oC was compressed from 300 mL to 0.360 mL and its  new pressure was found to be 400.0 kPa. What was the original pressure in  kPa?

1
Expert's answer
2021-01-12T07:04:09-0500

T = 20 ºC = constant

"V_1 = 300 \;mL \\

V_2 = 0.360 \;mL \\

p_2 = 400.0 \;kPa \\

p_1V_1 = p_2V_2 \\

p_1 = \frac{p_2V_2}{V_1} \\

= \frac{400.0 \times 0.360}{300.0} \\

= 0.48 \;kPa"

Answer: 0.48 kPa